1.) Helena earns $2000 this year and will earn $1500 the next year. The interest rate between each year is 10%. She wants to consume the same quantity each year. She wants to have spent all her money by the end of year 2.
Warning: each question does build on each other
a.) How much would she save? How much would she consume?
b.) If she wins the lottery during the first year, raising her earnings to $10000 this year, how much would she save, how much would she consume?
c.) If she now wants to consume twice as much in year 1 as she does in year 2, how much would she save, how much would she consume?
If she now has no preferences about her consumption, but she wants to save $5000 in year 1 (to buy a new car in year 2), how much would she consume in year 1 and year 2?
Answer #1

solution:- A Given Income of present yeaut= $2000. let us assume amount used for Consumption = x > Amount saved (Poesent yearB In the same u way, for current year income = 10,000 X= 1500+ C10,000->) + 10) () 9,000-)) x = 1500+ li(10,000- xtl•lx= 125